# The Causal Effect of the Two-For-One Strategy in the National Basketball Association

### v1

@@ 3 Methodology
Using the definitions from the previous section, we define the treatment $W$ for observation $i$ as $W_{i}=1$ if the TFO-O turned into a TFO-A and $W_{i}=0$ if the TFO-O turned into a TFO-NA. We let $Y_{i}$ be the POD. We employ the potential outcomes framework in order to define the primary causal effect of interest. Let $Y_{i}(1)$, $Y_{i}(0)$ be the potential outcomes under treatment and control, respectively. That is, $Y_{i}(1)$ is the POD a team would get if it turns a TFO-O into a TFO-A, and $Y_{i}(0)$ is the POD a team would get if it turns a TFO-O into a TFO-NA. A natural estimand of interest is the average treatment effect (ATE):
@@
Other estimands may also be of interest for particular analyses. In particular, we can restrict the population of interest to be either the treatment or control group, leading to the average treatment effect on the treated (ATT),
$ATT=E[Y_{i}(1)-Y_{i}(0)|W_{i}=1],$ (3)
and the average treatment effect on the control (ATC),
$ATC=E[Y_{i}(1)-Y_{i}(0)|W_{i}=0].$ (4)
We will use $\tau$ throughout the paper to represent these treatment effects, with the specific type of effect given by context.
Unfortunately, we do not observe both $Y_{i}(1)$ and $Y_{i}(0)$ for any particular observation. For each TFO-A and TFO-NA, it is impossible to know what would have happened if the team had chosen a different path. This is known as the fundamental problem of causal inference. If the scenarios where a team chose a TFO-O were the same, on average, in all meaningful ways to the scenarios where a team chose a TFO-A, an assumption known as exchangeability, then we could estimate the ATE from the data as $\hat{E}(Y_{i}|W_{i}=1)-\hat{E}(Y_{i}|W_{i}=0)$. Unfortunately that assumption is not reasonable in this situation, as potentially confounding variables such as the score margin and the time left in the quarter are likely to influence a team’s decision.
The methods of causal inference have been developed for precisely such circumstances, when a causal effect estimate is desired for observational data where exchangeability does not hold. These methods typically rely on a common set of assumptions, namely consistency, no interference, conditional exchangeability, and positivity. Below we discuss these assumptions and why they are reasonably satisfied for our data, followed by a description of the specific methods we will be using to estimate treatment effects.
@@ 3.1 Causal Assumptions
The consistency assumption implies that there is a well-defined treatment with no hidden variations. In our setting we recognize that not all TFO-A look the same, as highlighted in the previous section. Some end in a shot while others end in a foul. In some cases the team that had the TFO-A gets the ball back, and in other cases it does not. However, because these differences are a part of the definition of our treatment, they are not hidden and do not form an egregious violation of the consistency assumption.
@@
where $W_{-i}$ represents the treatment vector excluding $i$. This assumption would be violated in our setting if the success of a particular team’s TFO-A was affected by whether another TFO-O in the study was an attempt or a non-attempt. It seems unlikely that this would be true if the two units are from separate teams. It is possible due to the season-long nature of the data that if a particular team has a large number of TFO-A relative to TFO-NA early in the year, the players may practice last-second shots more because they have been experienced them more in the course of play, and thus they could improve the outcome of the TFO-A later in the year. However, players already practice these types of shots because of their importance to game play, and it seems unlikely that the amount of time they spend practicing such shots substantially depends on how often they attempt a TFO strategy. Thus the no interference assumption seems to reasonably hold for our setting.
Conditional exchangeability implies that, given a set of covariates $X$, the treatment assignment $W$ is independent of the potential outcomes
@@
This allows us to assume that treated and untreated individuals with the same values of $X$ have comparable distributions of potential outcomes, thus enabling causal comparisons. Suppose our set of covariates consists of the score margin and the time left in the quarter. Then conditional exchangeability would say, for example, that TFO-O scenarios where a team is down by 4 with 40 seconds left in the quarter are the same, on average, for both the TFO-A and TFO-NA groups. The key to reasonably satisfying this assumption is choosing a set of covariates rich enough to include any major potential confounding variables.
The list of covariates we consider is given by Table 1. When considering whether teams convert a TFO-O into a TFO-A, perhaps the most obvious factor is how much time is left in the quarter. We also believe that there may be quarter specific effects and thus include indicators for the first three quarters. The scoring margin at the time of the TFO-O is included as a team may feel more of a need to maximize possessions if it is trailing in the game. The quality of the lineup on the floor for both teams is also important and is characterized by the different rating covariates. We use the betting spread to account for how close a game is expected to be, and the betting total to account for how high or low scoring the game is expected to be. We believe these are the most important sources of potential confounding and thus that conditional exchangeability is reasonably satisfied. It is also common to check this condition by looking at the balance of the covariates in the groups created by the methods discussed below, which we do in Section 4.1.
@@
The final assumption necessary is the positivity assumption, which says that each TFO-O must have a positive probability, conditional on our set of covariates, of being in either the treatment or control group i.e.,
$10<P(W_{i}=w_{i}\mid X_{i})<1,\quad w_{i}=0,$ 1 (7)
@@ 3.2 Causal Methods for Average Treatment Effects
To estimate treatment effects we make use of two common causal inference methods: inverse probability weighting (IPW) and matching. Both of these methods rely on the concept of a propensity score (Rosenbaum & Rubin, 1983), which is defined as the probability of receiving treatment given baseline covariates. For a binary treatment $W\in\{0,1\}$, the propensity score denoted by $e(X)$ is given by
@@
This scalar summary of multiple covariates facilitates covariate balance between treated and control groups, aiming to replicate the distributional balance of a randomized experiment. Propensity scores are generally unknown quantities but can be estimated using logistic regression or machine learning methods such as random forests or gradient boosting (Hastie et al., 2009), which can handle complex relationships between covariates and treatment assignment (Lee et al., 2010). For our analysis we employ the covariate balancing propensity score method of Imai & Ratkovic (2013), which improves performance by leveraging both the goal of modeling treatment assignment and that of balancing the covariates. In practice, propensity scores are often computed in software such as R, using packages like MatchIt, which provides matching methods, or WeightIt, which computes weights for propensity score weighting (Stuart et al., 2011; Greifer, 2024b). Under the assumptions of conditional exchangeability, positivity and consistency, the propensity score allows for unbiased estimation of the ATE (Imbens & Wooldridge, 2009; Stuart, 2010).
Propensity scores are used to estimate treatment effects in several ways. IPW assigns weights to individuals to create a pseudo-population where the distribution of covariates is balanced between treatment groups. For an individual $i$ with covariates $X_{i}$ when the ATE is the estimand of interest, weights are computed as
$\frac{W_{i}}{\hat{e}(X_{i})}+\frac{1-W_{i}}{1-\hat{e}(X_{i})}$ (9)
@@
where $\hat{e}(X_{i})$ is the estimated propensity score. This weight is larger for individuals in groups with a lower probability of treatment assignment, adjusting for selection bias by up-weighting underrepresented individuals (Hirano et al., 2003). If conditional exchangeability holds in the true population, then marginal exchangeability holds in the pseudo-population. Thus the ATE can then be estimated by the weighted difference in outcomes:
$\hat{\tau}_{ATE}=\frac{1}{n}\sum_{i=1}^{n}\Big{(}\frac{W_{i}Y_{i}}{\hat{e}(X_{i})}-\frac{(1-W_{i})Y_{i}}{1-\hat{e}(X_{i})}\Big{)}$ (10)
where $Y_{i}$ denotes the observed outcome. If the ATT or ATC are the estimands of interest, the weights are simply modified so that the distribution of the group of interest is unchanged, leading to
$\hat{\tau}_{ATT}=\frac{1}{n}\sum_{i=1}^{n}\Big{(}W_{i}Y_{i}-\frac{\hat{e}(X_{i})(1-W_{i})Y_{i}}{1-\hat{e}(X_{i})}\Big{)},$ (11)
$\hat{\tau}_{ATC}=\frac{1}{n}\sum_{i=1}^{n}\Big{(}\frac{(1-\hat{e}(X_{i}))W_{i}Y_{i}}{\hat{e}(X_{i})}-(1-W_{i})Y_{i}\Big{)}.$ (12)
This inverse probability weighting (IPW) estimator can be implemented in R using the survey package, which enables the application of IPW techniques to survey data structures, and twang, which supports estimation and diagnostics for IPW models with large datasets (Robins et al., 2000; Cefalu et al., 2024)
Propensity score matching pairs treated and control individuals with similar propensity scores, aiming to create a balanced sample for which treatment and control groups are comparable on observed covariates. When estimating the ATT, each treated unit $i$ is matched to a control unit $j$ with a propensity score as close as possible to $e(X_{i})$. When estimating the ATC, each control unit $i$ is matched to a treated unit $j$ with a propensity score as close as possible to $e(X_{i})$. Matching can be done using nearest-neighbor matching, which minimizes the absolute difference between scores; caliper matching, which restricts matches to pairs within a predefined range; or optimal matching, which minimizes the overall distance across matched pairs (Austin, 2011; Rosenbaum, 2002). If there are many more observations in one of the groups, several units in the larger group may be matched with a single unit in the smaller group. Typically matching is done without replacement, although with replacement matching is also possible.
Again, if conditional exchangeability holds in the true population, then marginal exchangeability holds in the matched population, and the treatment effect can be estimated as the difference in the treatment and control group averages in the matched population. Thus for the ATT we have
$\hat{\tau}_{ATT}=\frac{1}{n_{1}}\sum_{i:W_{i}=1}\Big{(}Y_{i}-\frac{1}{|\mathscr{M}(i)|}\sum_{j\in\mathscr{M}(i)}Y_{j}\Big{)},$ (13)
where $n_{1}$ is the number of treated individuals and $\mathscr{M}(i)$ represents the matched control units for treated unit $i$. For the ATC we can simply swap the treatment and control group in the above formula. This method can effectively balance covariates if quality matches can be found, making it a robust approach for estimating causal effects when assumptions hold. In R, the MatchIt package provides flexible options for implementing nearest-neighbor, caliper, and optimal matching, along with diagnostic tools for balance checking (Stuart et al., 2011).
It is important to note the methodological distinction between these two methods. In IPW, each observation is reweighted based on the estimated propensity score, allowing us to balance the treated and control groups in expectation. In contrast, propensity score matching directly pairs treated units with control units that have similar propensity scores, creating a matched sample for estimation. Both approaches aim to reduce bias in estimating the treatment effect, but matching emphasizes the creation of comparable subgroups, whereas weighting adjusts the contribution of all observations based on their propensity scores. Both methods are widely used in causal inference applications. See Hernan & Robins (2024) for a theoretical account of both methods, and Shiba & Kawahara (2021) for a recent practical comparison. We present effect estimates from both methods in Section 4.1.
@@ 3.3 Causal Methods for Heterogenous Treatment Effects
The ATE provides insight into the overall causal effect for a population. However, treatment effects often vary across individuals within the population. The presence of heterogeneity in treatment effects can obscure important subpopulation dynamics if only the ATE is considered. Identifying and analyzing these differences allows for more tailored interventions, improving both individual and aggregate outcomes. Furthermore, examining treatment effect heterogeneity can enhance our understanding of how covariates interact with the treatment, providing deeper insights into causal mechanisms. This motivates the need for methodologies that move beyond the ATE to uncover nuanced variations in causal effects across different subgroups.
The Conditional Average Treatment Effect (CATE) (Athey & Imbens, 2016) is a statistical measure used to capture the heterogeneity in treatment effects across different subgroups within a population. It extends the concept of the ATE, which calculates the expected difference in outcomes due to a treatment across the entire population, by conditioning on specific covariates $X$. This approach is particularly useful in causal inference to identify how treatment efficacy varies across different individuals or groups, allowing for more personalized interventions.
@@
Classical approaches to heterogeneous treatment effect (HTE) estimation often rely on subgroup analysis, where subgroups are defined by observed covariates. However, traditional regression-based methods may be limited by the need to pre-specify interactions or assume homogeneity within subgroups. Recent advances have thus focused on machine learning techniques, which allow for more flexible, data-driven discovery of treatment effect heterogeneity. For instance, Athey & Imbens (2016) introduced causal trees for identifying subpopulations with distinct treatment effects, while Wager & Athey (2018) extended this to causal forests, a method based on random forests that provides both individual-level treatment effect estimates and uncertainty quantification.
Propensity score-based methods are also adapted for HTE analysis, such as subclassification on propensity scores within stratified groups (Rosenbaum & Rubin, 1983). However, propensity-based models assume covariate balance is achieved within each stratum, which may not hold in practice for high-dimensional data. To address this, flexible machine learning models, including generalized additive models and Bayesian Additive Regression Trees (BART), are commonly used to improve balance across strata and yield robust HTE estimates (Hill, 2011).
In our analysis, we use the causal forest framework to compute heterogenous treatment effects.
3.3.1 Causal Forest
Causal forests, introduced by Athey et al. (2019), Athey & Wager (2019), and Wager & Athey (2018), build on the framework of random forests (Breiman, 2001; Hastie et al., 2009) to estimate the conditional average treatment effect (CATE), denoted as $\tau(X_{i})$, rather than the outcome $Y_{i}$. These forests have become a powerful tool for causal inference, leveraging machine learning techniques to uncover treatment effect heterogeneity. The methods are implemented in the R package grf (Tibshirani et al., 2024), which serves as a comprehensive toolkit for causal forest analysis. Below, we detail the core components of this algorithm.
Orthogonalization via Double Machine Learning
A central challenge in causal inference is ensuring unbiased estimation of $\tau(X_{i})$ in the presence of confounding variables, as we have discussed earlier. Causal forests address this through an orthogonalization procedure inspired by the double machine learning framework of Chernozhukov, Chetverikov, Demirer, Duflo, Hansen, Newey & Robins (2018). This process involves ”residualizing” the outcome and treatment to remove the influence of confounders.
Given a dataset ${(Y_{i},W_{i},X_{i}),i=1,\dots,n}$, the first step is to estimate the expected outcome, $\mu(X_{i})$, for both treated and control groups. This outcome model predicts the conditional mean:
$\hat{\mu}(X_{i})=\mathbb{E}[Y_{i}|X_{i}=x,W_{i}=w]$
Next, the propensity score model estimates $e(X_{i})$, the probability of receiving treatment given the covariates:
$\hat{e}(X_{i})=\mathbb{P}(W_{i}=1|X_{i}=x)$
Using these models, the observed outcomes and treatment are residualized:
$\tilde{Y_{i}}=Y_{i}-\hat{\mu}(X_{i}),\quad\tilde{W_{i}}=W_{i}-\hat{e}(X_{i})$
Residualization ensures that the residualized outcome $\tilde{Y_{i}}$ is orthogonal to the covariates $X_{i}$, effectively isolating the treatment effect. These residuals serve as inputs for the subsequent tree-building process, reducing bias and enhancing the robustness of the treatment effect estimation.
Splitting Criterion: Maximizing Treatment Effect Heterogeneity
Causal forests partition the covariate space $\mathscr{X}$ by recursively splitting it to maximize treatment effect heterogeneity. Unlike regression trees, which minimize the variance of the outcome, causal trees aim to identify regions where treatment effects differ the most. For a split on a covariate $X_{j}$ at a threshold $s$, the data is divided into two child nodes:
$L=\{i:X_{i,j}\leq s\}\quad R=\{i:X_{i,j}>s\}$
The splitting criterion evaluates the difference in CATE between these nodes, maximizing the squared difference:
$\Delta(s)=\Big{(}\frac{1}{|L|}\sum_{i\in L}\hat{\tau}(X_{i})-\frac{1}{|R|}\sum_{i\in R}\hat{\tau}(X_{i})\Big{)}^{2}$
This approach encourages the tree to uncover splits that highlight heterogeneity in treatment responses, tailoring the model to detect meaningful differences across subpopulations.
Honest Estimation and Inference
To prevent overfitting and ensure valid inference, causal forests employ an honest estimation strategy, as proposed by Athey & Wager (2019). This involves splitting the data into two subsets: one for constructing the tree structure (splitting sample) and another for estimating treatment effects (estimation sample). For a given tree structure $T$, the treatment effect for each leaf $l\in T$ is estimated using:
$\hat{\tau}(l)=\frac{1}{|l|}\sum_{i\in l}\frac{\tilde{Y_{i}}\tilde{W_{i}}}{\hat{e}(X_{i})(1-\hat{e}(X_{i}))}$
Honest splitting mitigates the risk of overfitting by ensuring that data used to define the tree splits are separate from those used for effect estimation, thereby producing unbiased treatment effect estimates.
Ensemble of Trees: Aggregation for Stability
Causal forests aggregate multiple trees, each grown on a different bootstrap sample of the data, to enhance stability and reduce variance. For an individual $i$, the final estimate of the CATE is obtained by averaging the treatment effects from all trees in the ensemble:
$\hat{\tau}(X_{i})=\frac{1}{B}\sum_{b=1}^{B}\hat{\tau}_{b}(X_{i})$
This ensemble approach capitalizes on the diversity of tree structures to produce robust and reliable estimates of treatment effects.
In summary, causal forests combine orthogonalization, heterogeneity-driven splits, honest estimation, and ensemble averaging to provide a powerful framework for estimating treatment effect heterogeneity. This methodological rigor ensures that causal forests remain at the forefront of modern causal inference.
3.4 Assessing Heterogeneity
Once the causal forest model is built, it is essential to test for significant heterogeneity in the estimated causal effects. One approach to evaluate this is through a Calibration Test (Chernozhukov, Demirer, Duflo & Fernandez-Val, 2018), which is conducted as follows:
We fit a linear regression where the target (outcome) variable serves as the dependent variable, and the independent variables are: (1) the overall treatment effect, denoted as mean forest prediction, and (2) the predicted conditional average treatment effect (CATE) based on the out-of-bag trees, denoted as differential forest prediction.
If significant heterogeneity exists in the causal effects, the coefficient associated with differential forest prediction will be statistically different from zero. This indicates that the predicted heterogeneity in treatment effects contributes meaningfully to explaining the outcome variability.
The Rank Weighted Average Treatment Effect (RATE) (Yadlowsky et al., 2024) provides a more formal approach to testing for heterogeneity in causal effects. This method begins with a prioritization rule based on covariates, $X_{i}$. The prioritization rule, $S(X_{i})$, assigns a score to each sample, where a higher score reflects a higher expected benefit from the treatment. In other words, samples predicted to gain more from the treatment receive higher prioritization scores.
The method then measures the incremental benefit of administering the treatment to the top $q$of samples according to their prioritization scores, a concept captured by the Target Operating Curve (TOC). Let $F$ be the distribution function of $S(X_{i})$ and for $q\in(0,1]$, the TOC is defined as:
$\text{TOC}(q)=E[Y_{i}(1)-Y_{i}(0)\mid S(X_{i})\geq F^{-1}_{S(X_{i})}(1-q)]-E[Y_{i}(1)-Y_{i}(0)]$ (15)
At $\text{TOC}(1)$, the difference between the two expectations becomes zero, as the entire sample is included. The $\text{TOC}(q)$ is estimated by comparing the overall average treatment effect (ATE) with the ATE for the subset of samples whose prioritization scores exceed a given threshold, analogous to the receiver operating characteristic (ROC) curve.
RATE is the area under the curve of the TOC curve i.e.,
[…]
The RATE value lies between 0 and 1. A good prioritization rule, one that accurately reflects treatment benefit, will result in a RATE value close to 1. Conversely, a poor rule will yield a RATE value near 0.
To test for overall heterogeneity, we can use the CATE estimates as the prioritization rule, assigning higher scores to observations with higher predicted CATE values. If significant heterogeneity is present, the RATE will be close to 1.
Alternatively, we can assess heterogeneity across a specific variable by basing the prioritization rule on that variable rather than the CATE estimates. For instance, we could use $S(X_{i})=X_{i1}$ to rank samples by a particular covariate. In our case, we are interested in evaluating treatment effect heterogeneity when ranking is based on player rating and opponent player rating.
We now apply the methods discussed in the previous section to our data and present the results in the following section.

### v3

@@ 3 Methodology
Using the definitions from the previous section, we define the treatment $W$ for TFO opportunity $i$ as $W_{i}=1$ if the TFO is attempted and $W_{i}=0$ if the TFO is not attempted. We let $Y_{i}$ be the POD. We employ the potential outcomes framework (Splawa-Neyman, 1923; Rubin, 1974) in order to define the primary causal effect of interest. Let $Y_{i}(1)$, $Y_{i}(0)$ be the potential outcomes under treatment and control, respectively. That is, $Y_{i}(1)$ is the POD a team would get if it attempts a TFO at opportunity $i$, and $Y_{i}(0)$ is the POD a team would get if it does not attempt a TFO at opportunity $i$. A natural estimand of interest is the average treatment effect (ATE):
@@
The ATE quantifies how much larger the POD would be, on average, if NBA teams presented with a TFO opportunity always made a TFO attempt rather than never made a TFO attempt.
The fundamental problem of causal inference is that we do not observe both $Y_{i}(1)$ and $Y_{i}(0)$ for any particular observation. If the scenarios where a team attempted a TFO were the same, on average, in all meaningful ways to the scenarios where a team did not attempt a TFO, an assumption known as exchangeability, then we could estimate the ATE from the data as $\hat{E}(Y_{i}|W_{i}=1)-\hat{E}(Y_{i}|W_{i}=0)$. Unfortunately that assumption is not reasonable in this situation, as potentially confounding variables such as the score margin and the time left in the period are likely to influence a team’s decision.
The methods of causal inference have been developed for precisely such circumstances, when a causal effect estimate is desired for observational data where exchangeability does not hold. Many of these methods rely on a common set of assumptions, namely consistency, no interference, conditional exchangeability, and positivity. Below we discuss these assumptions and why they are reasonably satisfied for our data, followed by a description of the specific methods we will be using to estimate treatment effects.
@@ 3.1 Causal Assumptions
The consistency assumption implies that there is a well-defined treatment with no hidden variations. In our setting we recognize that not all TFO attempts look the same, as highlighted in the previous section. Some end in a shot while others end in a foul. In some cases the team that had the TFO attempt gets the ball back, and in other cases it does not. However, because these differences are a part of the definition of our treatment, they are not hidden and do not form an egregious violation of the consistency assumption.
@@
where $W_{-i}$ represents the treatment vector excluding $i$. This assumption would be violated in our setting if the success of a particular team’s TFO attempt was affected by whether another TFO opportunity in the study was an attempt or a non-attempt. One potential source of interference could come from the possibility of multiple TFO opportunities in the same period. For example, suppose a team gains possession with 43 seconds left and makes a shot with 36 seconds left, after which the opposing team gains possession and makes a shot with 29 seconds left. Then both teams have attempted a TFO. However, we do not believe that the success of the TFO attempt by the team gaining possession with 36 seconds left in this scenario depends on the fact that the first team is also attempting a TFO. Had the first team gained possession with 50 seconds left instead of 43, we would not consider this a TFO attempt but the second team gets the ball in an practically identical situation as the original scenario. Thus we have left these instances in our data set.
A slightly more complicated scenario occurs if the same team has back-to-back TFO opportunities. This may happen, for example, if a team gains possession with 43 seconds left, misses a shot with 36 seconds left, but gets an offensive rebound, thus earning another possession. Often in this case the player that gets the rebound is able to take a quick shot close to the goal which has a higher than average probability of going in or drawing a foul. So it is plausible that these TFO attempts may lead to a higher POD because they follow another TFO attempt from the same team. Thus we have removed these instances from our data set.
Other than units occuring in the same period, it is difficult to imagine a scenario where there would be a serious violation of this assumption. It is possible due to the season-long nature of the data that if a particular team has a large number of TFO attempts relative to non-attempts early in the year, the players may practice last-second shots more because they have been experienced them more in the course of play, and thus they could improve the outcome of the TFO attempts later in the year. However, players already practice these types of shots because of their importance to game play, and it seems unlikely that the amount of time they spend practicing such shots substantially depends on how often they attempt a TFO strategy. Thus the no interference assumption seems to reasonably hold for our setting.
Conditional exchangeability implies that, given a set of covariates $X$, the treatment assignment $W$ is independent of the potential outcomes:
@@
This allows us to assume that treated and untreated individuals with the same values of $X$ have comparable distributions of potential outcomes, thus enabling causal comparisons. Suppose our set of covariates consists of the score margin and the time left in the period. Then conditional exchangeability would say, for example, that TFO opportunities where a team is down by 4 with 40 seconds left in the period are the same, on average, for both the set of attempts and non-attempts.
One way to examine the reasonableness of this assumption is to model the relationships between treatment, response, and covariates in a directed acyclic graph (DAG). In these diagrams, an arrow from one node to another represents a causal relationship between those variables. For a DAG to be complete, it must contain all common causes of any variables in the graph. Under this assumption, we can use a quick visual check called the back-door criteria to determine whether a set of covariates satisfies the conditional exchangeability assumption. See Pearl (2009) for more information on the use of DAGs in causal inference.
We propose the DAG given in Figure 2 as a model for our causal structure, using the covariates in Table 1. Note that we combine the ratings variables into a single node for the sake of simplicity and because we think they would have the same relationship structure with the other variables. We think that each of these covariates could reasonably affect both the treatment and response variables. Teams may be more likely to attempt a TFO if there is more time left in the period, and this also could possibly allow for more possessions after the TFO opportunity, which could increase the POD. There spread and the score margin represent the expected and current point difference between the teams, and it is possible that teams that are behind or are expected to be behind may feel more of a need to maximize possessions. At the same time these teams may score less on average per possession than those winning or expected to win, affecting the POD. The total score is in some ways a proxy for pace of play. Teams that play faster may be more likely to attempt a TFO than a more methodical team, and a faster pace can lead to more possessions and thus affect the POD. The quality of the players on the floor certainly impacts the scoring at the end of the period, but it may also impact the likelihood of a TFO attempt if, for example, a team playing its best players is able to execute well against an opponent playing weaker players. Finally, we allow for the existence of period specific effects as team strategies may well be different during different parts of the game.
Figure 2: DAG of our proposed causal structure, created using the DAGitty web application (Textor et al., 2017)
In addition to these confounding relationships, there are several relationships between covariates included in the DAG. For example, some coaches have set substitution patterns, and so the period of the game affects the ratings of the players on the floor. Alternatively, if a team is ahead by a large amount, a coach may change this pattern and rest his best players more than in a close game. It is important that none of these extra relationships lead to a cycle in our graph, which is not allowed in a DAG and would violate temporal restrictions of causality. Letting $X$ be the set of covariates listed in our DAG, the back-door criteria and thus conditional exchangeability is satisfied. However, while we believe that we have included the most important sources of confounding, there is the possibility of unmeasured confounding in any causal analysis. In Section 4.1.3 we examine the sensitivity of our estimates to unmeasured confounding. It is also appropriate to check conditional exchangeability by looking at the balance of the covariates in the groups created by the methods discussed below, which we do in Section 4.1.1.
@@
The final assumption necessary is the positivity assumption, which says that each TFO opportunity must have a positive probability, conditional on our set of covariates, of being in either the treatment or control group i.e.,
$0<P(W_{i}=w_{i}\mid X_{i})<1,\quad w_{i}=0,1$ (4)
@@ 3.2 Causal Methods for Average Treatment Effects
Many causal inference methods rely on the concept of a propensity score (Rosenbaum and Rubin, 1983), which is defined as the probability of receiving treatment given baseline covariates $X$. For a binary treatment $W\in\{0,1\}$, the propensity score denoted by $e(X)$ is given by
@@
Propensity scores are generally unknown quantities but can be estimated using logistic regression or machine learning methods.
@@
Propensity scores are used to estimate treatment effects in several ways. Inverse probability weighting (IPW) is a class of estimators that assigns weights to individuals to create a pseudo-population where the distribution of covariates is balanced between treatment groups. These weights are larger for individuals in groups with a lower probability of treatment assignment, adjusting for selection bias by up-weighting underrepresented individuals (Hirano et al., 2003). If conditional exchangeability holds in the true population, then marginal exchangeability holds in the pseudo-population. Thus the ATE can then be estimated by the weighted difference in outcomes,
$\hat{ATE}_{IPW}=\frac{1}{n}\sum_{i=1}^{n}\Big(\frac{W_{i}Y_{i}}{\hat{e}(X_{i})}-\frac{(1-W_{i})Y_{i}}{1-\hat{e}(X_{i})}\Big),$ (6)
where $\hat{e}(X_{i})$ is the estimated propensity score.
Note that the set of covariates $X_{i}$ contains information about both the treatment assignment and the outcome. This motivates an alternative approach known as augmented inverse probability weighting (AIPW) (Glynn and Quinn, 2010). In addition to estimating propensity scores, this method estimates the outcome under both treatment and control, denoted $\hat{E}(Y_{i}|W_{i}=1,X_{i})$ and $\hat{E}(Y_{i}|W_{i}=0,X_{i})$, respectively. The IPW estimator is then adjusted by a term involving the weighted average of these two estimators:
$\hat{ATE}_{AIPW}=\frac{1}{n}\sum_{i=1}^{n}\left\{\Big(\frac{W_{i}Y_{i}}{\hat{e}(X_{i})}-\frac{(1-W_{i})Y_{i}}{1-\hat{e}(X_{i})}\Big)-\frac{W_{i}-\hat{e}(X_{i})}{\hat{e}(X_{i})(1-\hat{e}(X_{i}))}\right.\\ \left.[1-\hat{e}(X_{i})\hat{E}(Y_{i}|W_{i}=1,X_{i})+\hat{e}(X_{i})\hat{E}(Y_{i}|W_{i}=0,X_{i})]\right\}.$ (7)
This AIPW estimator is doubly-robust in that it is consistent for the ATE whenever either the propensity score model or the outcome model is correctly specified. It also requires the conditional exchangeability assumption mentioned above.
@@ 3.3 Causal Methods for Heterogenous Treatment Effects
While the ATE summarizes the overall impact of executing a TFO, it is natural to expect that its effect varies across different possession contexts. For example, some teams are naturally more effective at generating quick, quality scoring chances under time pressure, whereas others tend to rely on slower, more structured offensive sets that may yield lower-quality shots when rushed. Likewise, a TFO taken when a team holds a comfortable lead may influence subsequent decision-making differently than one taken in a highly competitive situation. The defensive characteristics of the opponent also matter: facing a unit that is slower to adjust to rapid possessions may create more favorable scoring conditions than facing a unit that is disciplined and well-positioned. These contextual differences may naturally lead to heterogeneous treatment effects. Using causal forests to detect this heterogeneity helps clarify when and in what types of game states the 2-for-1 is most advantageous, complementing the ATE without affecting its interpretation or underlying assumptions.
The Conditional Average Treatment Effect (CATE) (Athey and Imbens, 2016) is a statistical measure used to capture the heterogeneity in treatment effects across different subgroups within a population. It extends the concept of the ATE, which calculates the expected difference in outcomes due to a treatment across the entire population, by conditioning on specific covariates $X$. This approach is particularly useful in causal inference to identify how treatment efficacy varies across different individuals or groups, allowing for more personalized interventions.
@@
Classical approaches to heterogeneous treatment effect (HTE) estimation typically rely on subgroup analysis defined by observed covariates. Regression-based methods require pre-specifying interactions or assume homogeneity within subgroups, which can limit their ability to capture complex patterns. Recent work therefore focuses on flexible machine learning methods that discover heterogeneity in a data-driven manner. For example, Athey and Imbens (2016) introduced causal trees to identify subpopulations with different treatment effects, and Wager and Athey (2018) extended this framework to causal forests, providing individual-level treatment effect estimates along with measures of uncertainty. Propensity score–based approaches Rosenbaum and Rubin (1983) can also be adapted for HTE analysis, though subclassification may struggle in high dimensions; methods such as BART (Hill, 2011) and other flexible learners are commonly used to mitigate these issues. In our analysis, we estimate HTEs using causal forests.
3.3.1 Causal Forests
Causal forests (Athey et al., 2019; Athey and Wager, 2019; Wager and Athey, 2018) are flexible, nonparametric methods for estimating conditional average treatment effects (CATEs) and exploring treatment effect heterogeneity across covariates. They extend the random forest framework (Breiman, 2001) by adapting both the splitting rules and estimation strategies to the causal inference setting. The procedure begins with an orthogonalization step, inspired by the double machine learning framework (Chernozhukov et al., 2018) in which both the outcome and treatment variables are residualized with respect to observed covariates. This removes variation explained by potential confounders, ensuring that differences in estimated treatment effects are not driven by baseline imbalances. Residualized variables are then used to grow decision trees where the splitting criterion is chosen to maximize heterogeneity in treatment effects between child nodes, rather than minimizing prediction error of the outcome. This design allows the trees to uncover complex, nonlinear patterns in how treatment effects vary across the population.
To prevent overfitting and obtain valid inference, causal forests employ an “honest” estimation strategy: the data are split into two parts, with one subsample used to determine the structure of the tree (i.e., the splits) and the other used to estimate treatment effects within the resulting leaves. This honest splitting ensures that the estimated treatment effects are asymptotically unbiased. Many such trees are grown on bootstrap samples of the data, and their predictions are aggregated to form the final CATE estimate for each observation, reducing variance and improving stability. Because orthogonalization reduces confounding and the splitting rule targets heterogeneity, causal forests can uncover treatment effect patterns that parametric models may miss. They also handle high-dimensional covariates and complex interactions while providing valid inference. These properties make causal forests well-suited for identifying subpopulations with differing treatment responses.